応用流体力学特論 2 単位 (選択)

Applied Fluid Dynamics  石原国彦 · 教授/ 知的力学システム工学専攻機械創造システム工学コース機械システム講座

武藤裕則 · 教授/ 知的力学システム工学専攻建設創造システム工学コース環境整備工学講座

【授業目的】河川における洪水流と流砂の諸特性を解説したのち，それらの数値 解析技術について基礎から応用までを習得させることを目的とする． 【授業概要】まず，河川流の解析に使用される基礎式，流送土砂量の特性式およ び河床変動の基礎式を解説する．ついで，偏微分方程式の形態を有する現象 の支配方程式を主として有限差分法によって数値解析する手法について講述 したのち，簡単な問題を与えて，解析プログラム (FORTRAN) 構築の実習を させる． 【授業形式】講義および演習 【キーワード】洪水流，流砂，河床変動，数値解法 【先行科目】 [先行科目] 【関連科目】『地域防災学特論』 (0.7)，『水循環工学特論』 (0.7) 【履修要件】学部において水工学，河川工学，数値解析などの科目を履修したう えでの受講が望ましい． 【履修上の注意】特になし． 【到達目標】河川の流れと河床変動の基本的性質を理解するとともに，それらの 実用的数値解法を修得している． 【授業計画】 1. ガイダンス (河道問題と数値解析) 2. 河川流の基礎方程式 3. 砂礫の掃流限界と浮遊限界 4. 掃流砂量式 5. 浮遊砂量式 6. 河床変動の基礎方程式 7. 理解度テスト (1) 8. 微分方程式の数値解法概説 9. 流れの数値解法 (1) 10. 流れの数値解法 (2) 11. 河床変動の数値解法 (1) 12. 河床変動の数値解法 (2) 13. プログラミング課題の出題と解説 14. プログラミング実習 (1) 15. プログラミング実習 (2)

16. 理解度テスト (2)· 課題発表【成績評価基準】理解度テスト (1) および (2) の成績を 6:4 の比率で総合化し， 60%以

上を合格とする．【教科書】 (社) 砂防学会編，山地河川における河床変動の数値計算法，山海堂．【参考書】特に指定しない．【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216579 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 岡部 .

Applied Fluid Dynamics 2 units (selection)

Kunihiko Ishihara · Professor / Mechanical Systems, Mechanical Engineering, Intelligent Structures and Mechanics Systems Engineering Yasunori Muto · Professor / Environmental Conservation Engineering, Civil and Environmental Engineering, Intelligent Structures and Mechanics Systems Engineering

Target⟩ This subject is concerned with flood flow, sediment transport and riverbed variation. The aim of this subject is to understand fundamentals of those phenomena as well as numerical solution methods for the basic equations.

Outline⟩ First, characteristics of flood flow, sediment transport and riverbed variation are interpreted and the basic equations of those phenomena are deduced and presented. Next, fundamentals of numerical solution methods for partial differential equations are lectured to be put into practical usage for numerical simulation of various kinds of flood flow and riverbed variation. Finally, some simple problems are given to practice computer programming by FORTRAN.

Style⟩ Lecture and excercise �������⟩ flood flow, sediment transport, riverbed variation, numerical model Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Advanced Disaster Reduction Engineering”(0.7), “Advanced

Water Circulation Engineering”(0.7) Requirement⟩ Fundamental knowlege of hydraulics, river engineering and

numerical analysis is necessary. Notice⟩ not specified. Goal⟩ fundamentals of flood flow, sediment transport and riverbed variation are

understood and numerical techniques for those problems are aquired at practical level.

Schedule⟩

1. guidance (river problems and numerical analysis)

2. basic equations of river flow

3. critical condition of sediment motion

4. formulas of bed load

5. formulas of susupended load

6. basic equations of riverbed variation

7. test(1)

8. outline of numerical solution of differential equations

9. numerical models for river flow (1)

10. numerical models for river flow (2)

11. numerical models for riverbed variation (1)

12. numerical models for riverbed variation (2)

13. problem setting and orientation for programming

14. practice of programing (1)

15. practice of programing (2)

16. test(2), presentation

Evaluation Criteria⟩ marks of tests (1) and (2) are summed with the weights of

6 and 4, respectively, and the passing mark is 60%. Textbook⟩ Japan Sabo-Gakkai, Numerical solution of riverbed variation in

mountain rivers, Sankaido Publ.. Reference⟩ not specified.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216579

Student⟩ only for specified course

Contact⟩ ⇒ Okabe .